Geological Formation Group Failed Total Gneiss 170 1885 Granite 2 28 Loch raven schist 523 3733 Mafic 14 363 Marble 29 309 PrPLEASE SHOW STEP BY STEP SOLUTION AND GIVE CORRECT ANSWER (VERY IMPORTANT).
Answer #1
failed total
170 1885
2 28
523 3733
14 363
29 309
60 1403
46 933
3 39
847 8693

a)

P(A and B) = (170 + 523 + 60 ) / 8693

= 753/8693 = 0.0866

b)

P(A or B) = P(A) + P(B) - P(A and B)

P(A) = (1885 + 3733 + 1403 )/ 8693 = 0.8077

P(B) = 847/8693 = 0.0974

hence

P(A or B) = 0.8077 + 0.0974 - 0.0866

= 0.8185

c)

= 1 - P(A and B)

= 1- 0.0866

= 0.9134

d)

P(A) = 0.8077

Please rate

Please look below in comments to get help in other questions